Prep corporation acquired the Snap company with 80% by paying cash 500,000 on 1 Jan 2009 Snap Net income during the year is 30,000 and declare dividend 15,000 The Excess amount is allocated to identifiable asset and liabilities Following is the Balance Sheet as at 1 Jan 2009 Asset Prep Snap Cash Account Recievable Other Current Asset Equipment, Net Building Net Land Total Assete 100,000 50,000 120,000 200,000 50,000 150,000 670,000 150,000 60,000 120,000 300,000 80,000 250,000 960,000 Liabilities Accoun Payable Note Payable Total Liabilities 250,000 150,000 400,000 300,000 250,000 550,000 Shareholder Equity Capital Retained Earning Total S.E 200,000 70,000 270,000 250,000 160,000 410,000 Please calculate 5- Dividend 6- What is the invenstment of Prep in Snap at the end 7- What is Non Controlling interest at the end of the year 8- Please prepare all the journal entries of the year
Answer #1

5. Calculation of dividend = 15000* 80% = 12000

6. Investment in Snap = Calculation of Intangible asset during acquisition = 625000(500000/80%) - 410000 = 215000

Total Asset as on Jan 1 2009 = 960000+215000 = 1175000

Calcuation of non controlling interest in Snap = 625000*20% = 125000

Calculation of Investment in Snap = 625000+15000 - 128000 = 512000

7. non controlling interest in Snap at the end of the year = 125000 + 20%*15000 = 128000

8. Proposed Dividend entry = Profit & loss a/c Dr. 15000

Proposed Dividend a/c cr. 15000

Dividend Paymen entry = Proposed Dividend a/c Dr. 15000

To cash a/c Cr. 15000

Transfer of Profit in P&L A/c

P&La/c Dr. 15000

To Retained Earning a/c Cr. 12000

To Non Controlling interest a/c Cr. 3000